OpenStudy (anonymous):

PLEASE HELP! Explain how applying the Associative Property of Addition to 7.5 + (2.5 + 4.9) makes it easier to simplify. Show the solution in your explanation.

OpenStudy (anonymous):

Add 7.5 nd 2.5 first

OpenStudy (skullpatrol):

The question asks you to explain how using the property: (a+b)+c=a+(b+c) will make your mental math easier. So, look at the numbers in 7.5 + ( 2.5 + 4.9 ) It is easier to add 7.5 + 2.5, FIRST to get 10.0 rather than adding 2.5 + 4.9 first as the question is written: 7.5 + ( 2.5 + 4.9 ) The Associative Property of Addition states that you can add whichever two numbers you want to do FIRST, and then add that sum to the third number: (a+b)+c=a+(b+c) Giving you 10.0 + 4.9 which is again easier to add because 10.0 has three zeros. The final answer being 14.9. Compare the above way to doing the question as it is written: 7.5 + ( 2.5 + 4.9 ) 7.5 + ( 7.4 ) = 14.9 will help you understand the associative property of addition.